OpenStudy (anonymous):

cot(-4pi/3) find the exact value of the following expression without using a calculator

OpenStudy (anonymous):

Use the unit circle.

OpenStudy (anonymous):

If it makes it easier -4pi/3 is equivalent to -240 degree

OpenStudy (amistre64):

trig is alot of memorizations ....

OpenStudy (anonymous):

im getting up to that, but then its saying that to find cot, i have to take cos/sin, and cos is square rt of 3 and when i do sin its -1/2 and the correct answer is suppose to be neg square rt of 3/3.

OpenStudy (amistre64):

the angle is whats important; all the trig functions are derived once you know the angle right?

OpenStudy (amistre64):

-4pi (180) ----* ----- = -4(60) = -240 3 pi

OpenStudy (amistre64):

360 - 240 = 120 degrees

OpenStudy (amistre64):

sin(120) = sin(60) = sqrt(3)/2 cos(120) = -cos(60) = -1/2 cot = cos/sin -1/2 -------- = -1/sqrt(3) sqrt(3/2)

OpenStudy (amistre64):

-1/sqrt(3) is a fine answer but alot of people believe that it is bad manners to have a sqrt(..) as a denominator so they rationalize it like this: -1 sqrt(3) -sqrt(3) ----- * ------ = ------- sqrt(3) sqrt(3) 3

OpenStudy (anonymous):

thank you, that explained it ten times better than the book ever could.

OpenStudy (anonymous):

while im on the subject, do you have any good math online resources? so far, i have my calc textbook, a calc book from kaplan, calc for dummies. plus youtube videos lol. its so miserable to learn math without a teacher.

OpenStudy (amistre64):

i taught meself calculus :) it is nice to have teacher to fill in the blanks yes. If its practice you need; interactmath.com has alot of free practice problems

OpenStudy (anonymous):

did you really? damn. mad props. some of it is easy, but some of it, forget about it. haha. thank youu